OpenStudy (anonymous):

how do you solve (1/2)(3pi/2)(6) how do you solve (1/2)(3pi/2)(6) @Mathematics

OpenStudy (amistre64):

hmmm .. id say multiply it out and simplify

OpenStudy (amistre64):

\[\frac{a}{b}*\frac{c}{d}=\frac{ac}{bd}\]

OpenStudy (anonymous):

Answer is 14.13 Do BEDMAS.